OpenStudy (anonymous):

If possible to factor the expression: x^8-1

OpenStudy (anonymous):

think about how to factor something of the form of a difference of squares, like a^2 - b^2

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

well, if you had x^2 - 1, could you factor it?

OpenStudy (anonymous):

Yes? Honestly, I don't know. Math isn't my subject at all. :(

OpenStudy (anonymous):

Have you learned this: a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

because when you multiply the right side out, you get a^2 + ab - ab - b^2

OpenStudy (anonymous):

so you can factor x^2 - 1 like this: (x + 1)(x - 1)

OpenStudy (anonymous):

because (x + 1)(x - 1) = x^2 + x - x - 1 = x^2 - 1

OpenStudy (anonymous):

so for your problem, you have x^8 - 1 x^8 is like x multiplied by itself 8 times.... (x)(x)(x)(x)(x)(x)(x)(x) so that is like (x^4)(x^4) which is (x^4) squared... So let's rewrite "x^8 - 1" as "x^4 squared minus 1" or "(x^4)^2 - 1" Now it's just like the form I showed you a minute ago... x^8 - 1 = (x^4 - 1)(x^4 + 1)

OpenStudy (anonymous):

Yeah, you can factor it.

OpenStudy (anonymous):

OHHHHHH MY GOODNESS!!!!! THAT MAKES SO MUCH MORE SENSE!!!!! Thank you!!! I think I'm catching on now!

OpenStudy (anonymous):

it is a little tricky... glad you are seeing it!!! :)

OpenStudy (anonymous):

Ok, I tried it, but the program I'm using said it was incorrect.